# Neural Eulerian Scene Flow Fields

### v2

@@ 5.3 Beyond Autonomous Vehicles
Due to a dearth of real-world, labeled scene flow data, prior scene flow work on real data overwhelmingly evaluates on autonomous vehicle datasets (Dewan et al., 2016; Li et al., 2021b; Jund et al., 2021; Li et al., 2023; Chodosh et al., 2023; Liu et al., 2024; Vedder et al., 2024; Khatri et al., 2024); consequently, motion understanding in other important domains like tabletop manipulation has been neglected. To showcase EulerFlow’s out-of-the-box flexibility and generalizability, we visualize EulerFlow on several dynamic tabletop scenes we collected using the ORBBEC Astra, a low cost depth camera commonly used in robotics (Figure 14). For viewing ease, we paint our point clouds with color; however, RGB information is not provided to EulerFlow during optimization. While EulerFlow only reasons about point clouds, it can leverage video mono depth estimates to describe RGB-only scene flow (Appendix A.2). Interactive visuals are available at vedder.io/eulerflow.
Figure 14: Visualizations of EulerFlow’s emergent 3D point tracking behavior that demonstrate the quality of its ODE estimate. Row 1 depicts tracking a tomato placed in the sink by a human hand; note the point does not move despite the hand grasping the tomato. Row 2 depicts tracking of painters tape rolling off a table; EulerFlow is able to estimate its trajectory even after it disappears out of frame. Row 3 depicts tracking of the motion of a jack commonly used in tabletop manipulation experiments (Venkatesh et al., 2023). Row 4 depicts tracking of a tennis ball taped to a flexible rod. All tracks are produced by Euler integration through the estimated ODE from the initial conditions shown in the left column. Note that point clouds are shown in color for visualization purposes only.
@@ 6.1 Limitations and Future Work
EulerFlow is expensive to optimize. With our implementation, optimizing EulerFlow for a single Argoverse 2 sequence takes 24 hours on one NVIDIA V100 16GB GPU, putting it on par with the original NeRF paper’s computation expense (Mildenhall et al., 2021). However, like with NeRF, we believe algorithmic, optimization, and engineering improvements can significantly reduce runtime.
EulerFlow does not understand ray casting geometry. During ego-motion, a static foreground occluding object casts a moving shadow on the background; this causes Chamfer Distance to estimate this as a leading edge of moving structure, encouraging false motion artifacts (Li et al., 2021b). This can be addressed with optimization losses that model point clouds as originating from a time of flight sensor with limited visibility, as has been successfully demonstrated in the reconstruction (Chodosh et al., 2024) and forecasting literature (Khurana et al., 2023; Agro et al., 2024), rather than an unstructured set of points to be associated via local point distance.
@@ A.1 How does the choice of learnable function class and design of encodings impact EulerFlow?
EulerFlow at its core is an optimization loop over a simple, feedforward ReLU-based multi-layer perception inherited from Neural Scene Flow Prior (Li et al., 2021b). How does this choice of learnable function class impact the performance of EulerFlow? To better understand these design choices we examine the choice of non-linearity and time feature encoding.
Figure 15: Mean Dynamic Normalized EPE of EulerFlow on the Argoverse 2 val split for less-smooth configurations of its learnable function class. These results indicate that the smoothness of the ReLU non-linearity proposed by Li et al. transfers well to EulerFlow.
One of Li et al.’s core theoretical contributions demonstrates that NSFP’s ReLU MLP is a good prior for scene flow because it represents a smooth learnable function class, and scene flow is often locally smooth with respect to input position. However, unlike NSFP, EulerFlow is fitting flow over a full ODE; while it seems reasonable to assume that this ODE is typically also locally smooth, cases like adjacent cars moving rapidly in opposite directions may benefit from the ability to model higher frequency, less locally smooth functions. To test this hypothesis, we ablate EulerFlow by replacing its normalized time with higher frequency sinusoidal time embeddings (mirroring Wang et al.’s proposed time embedding for NTP), as well as try other popular non-linearities like SinC (Ramasinghe et al., 2024) and Gaussian (Chng et al., 2022) from the coordinate network literature. Figure 15 features negative results on these ablations across the board; Gaussians were unable to converge due the extremely high frequency representation triggering early stopping, while the use of SinC and higher frequency time embeddings both resulted in worse overall performance, indicating that Li et al.’s smooth function prior does indeed seem appropriate for EulerFlow’s neural prior.
@@ A.2 EulerFlow with Monocular Depth Estimates
While EulerFlow only consumes point clouds, we can leverage RGB-based video monocular depth estimators to fit scene flow. In Figure 16, we use DepthCrafter (Hu et al., 2024) to generate a point cloud from the raw RGB of the tabletop video from Figure 14, Row 4.
Figure 16: Visualizations of EulerFlow’s emergent 3D point tracking behavior on monocular depth estimates from DepthCrafter (Hu et al., 2024). Interactive visualizations available at vedder.io/eulerflow.
@@ B.2 Why didn’t you use a Neural ODE or a Liquid Neural Network?
Neural ODEs (Chen et al., 2018) take variable size and number of steps in latent space to do inference; imagine a ResNet that can use an ODE solver to dynamically scale the impact of the residual block, as well as decide the number of residual blocks. They are not a function class specially designed to fit derivative estimates well. Similar to Neural ODEs, Liquid Neural Networks (Hasani et al., 2021) focus on the same class of problems and are similarly not applicable.
@@ B.3 Why didn’t you do experiments on FlyingThings3D / <simulated dataset>?
Most popular synthetic datasets do not contain long observation sequences (Mayer et al., 2016; Butler et al., 2012), but instead include standalone frame pairs. Our method leverages the long sequence of observations to refine our neural estimate of the true ODE. Indeed, on two frames, EulerFlow collapses to NSFP.
More importantly, these datasets are also not representative of real world environments. To quote Chodosh et al.: “[FlyingThings3D has] unrealistic rates of dynamic motion, unrealistic correspondences, and unrealistic sampling patterns. As a result, progress on these benchmarks is misleading and may cause researchers to focus on the wrong problems.” Khatri et al. also make this point by highlighting the importance of meaningfully breaking down the object distribution during evaluation identify performance on rare safety-critical categories. FlyingThings3D does not have meaningful semantics; it’s not obvious what things even matter or how to appropriately break down the scene.
@@ D.1 Formulating the ODE
Given a (possibly moving) particle in some canonical frame (i.e. time 00), we define a function $L(x_{0},y_{0},z_{0},t)$ that can describe its location at an arbitrary future time $t$, i.e. a Lagrangian description of motion (Figure 4).
@@ D.2 Arbitrary start and end times from the Eulerian formulation
In the above derivation, $L$ requires that a moving point be defined in terms of a canonical frame defined at time 00, as is common in the deformation in reconstruction literature. However, the Eulerian formulation has no such requirement, allowing us to select arbitrary start and end times across different point queries. To showcase this, we can query $F$ to extract the trajectory of a particle at $t$ across the range $[t,t^{\prime}]$ starting at $x_{t},y_{t},z_{t}$ simply by changing the range of the integral in Equation 10, i.e.
[…]
@@ D.3 Euler Integration to approximately solve the ODE
If $F$ is of arbitrary form and we want to compute the concrete values of $L$, we cannot exactly compute the continuous integral from 00 to $t$; we must approximate this with finite differences. Thus, we split the time range 00 to $t$ into $k$ steps, where each step is of size $\frac{t}{k}$. Thus, we can again define $L$ via recursion, but this time explicitly.

### v3

@@ 5.3 Beyond Autonomous Vehicles
Due to a dearth of real-world, labeled scene flow data, prior scene flow work on real data overwhelmingly evaluates on autonomous vehicle datasets (dewan2016rigid; nsfp; scalablesceneflow; fastnsf; chodosh2023; liu2024selfsupervisedmultiframeneuralscene; vedder2024zeroflow; khatri2024trackflow); consequently, motion understanding in other important domains like tabletop manipulation has been neglected. To showcase EulerFlow’s out-of-the-box flexibility and generalizability, we visualize EulerFlow on several dynamic tabletop scenes we collected using the ORBBEC Astra, a low cost depth camera commonly used in robotics (Figure 14). For viewing ease, we paint our point clouds with color; however, RGB information is not provided to EulerFlow during optimization. While EulerFlow only reasons about point clouds, it can leverage video mono depth estimates to describe RGB-only scene flow (Appendix A.2). Interactive visuals are available at vedder.io/eulerflow.
Figure 14: Visualizations of EulerFlow’s emergent 3D point tracking behavior that demonstrate the quality of its ODE estimate. Row 1 depicts tracking a tomato placed in the sink by a human hand; note the point does not move despite the hand grasping the tomato. Row 2 depicts tracking of painters tape rolling off a table; EulerFlow is able to estimate its trajectory even after it disappears out of frame. Row 3 depicts tracking of the motion of a jack commonly used in tabletop manipulation experiments (venkatesh2023samplingbased). Row 4 depicts tracking of a tennis ball taped to a flexible rod. All tracks are produced by Euler integration through the estimated ODE from the initial conditions shown in the left column. Note that point clouds are shown in color for visualization purposes only.
@@ 6.1 Limitations and Future Work
EulerFlow is expensive to optimize. With our implementation, optimizing EulerFlow for a single Argoverse 2 sequence takes 24 hours on one NVIDIA V100 16GB GPU, putting it on par with the original NeRF paper’s computation expense (nerf). However, like with NeRF, we believe algorithmic, optimization, and engineering improvements can significantly reduce runtime.
EulerFlow does not understand ray casting geometry. During ego-motion, a static foreground occluding object casts a moving shadow on the background; this causes Chamfer Distance to estimate this as a leading edge of moving structure, encouraging false motion artifacts (nsfp). This can be addressed with optimization losses that model point clouds as originating from a time of flight sensor with limited visibility, as has been successfully demonstrated in the reconstruction (chodosh2024simultaneousmapobjectreconstruction) and forecasting literature (khurana2023point; agro2024uno), rather than an unstructured set of points to be associated via local point distance.
@@ A.1 How does the choice of learnable function class and design of encodings impact EulerFlow?
EulerFlow at its core is an optimization loop over a simple, feedforward ReLU-based multi-layer perception inherited from Neural Scene Flow Prior (nsfp). How does this choice of learnable function class impact the performance of EulerFlow? To better understand these design choices we examine the choice of non-linearity and time feature encoding.
Figure 15: Mean Dynamic Normalized EPE of EulerFlow on the Argoverse 2 val split for less-smooth configurations of its learnable function class. These results indicate that the smoothness of the ReLU non-linearity proposed by nsfp transfers well to EulerFlow.
One of nsfp’s core theoretical contributions demonstrates that NSFP’s ReLU MLP is a good prior for scene flow because it represents a smooth learnable function class, and scene flow is often locally smooth with respect to input position. However, unlike NSFP, EulerFlow is fitting flow over a full ODE; while it seems reasonable to assume that this ODE is typically also locally smooth, cases like adjacent cars moving rapidly in opposite directions may benefit from the ability to model higher frequency, less locally smooth functions. To test this hypothesis, we ablate EulerFlow by replacing its normalized time with higher frequency sinusoidal time embeddings (mirroring ntp’s proposed time embedding for NTP), as well as try other popular non-linearities like SinC (ramasinghe2024on) and Gaussian (garf) from the coordinate network literature. Figure 15 features negative results on these ablations across the board; Gaussians were unable to converge due the extremely high frequency representation triggering early stopping, while the use of SinC and higher frequency time embeddings both resulted in worse overall performance, indicating that nsfp’s smooth function prior does indeed seem appropriate for EulerFlow’s neural prior.
@@ A.2 EulerFlow with Monocular Depth Estimates
While EulerFlow only consumes point clouds, we can leverage RGB-based video monocular depth estimators to fit scene flow. In Figure 16, we use DepthCrafter (hu2024DepthCrafter) to generate a point cloud from the raw RGB of the tabletop video from Figure 14, Row 4.
Figure 16: Visualizations of EulerFlow’s emergent 3D point tracking behavior on monocular depth estimates from DepthCrafter (hu2024DepthCrafter). Interactive visualizations available at vedder.io/eulerflow.
@@ B.2 Why didn’t you use a Neural ODE or a Liquid Neural Network?
Neural ODEs (neuralode) take variable size and number of steps in latent space to do inference; imagine a ResNet that can use an ODE solver to dynamically scale the impact of the residual block, as well as decide the number of residual blocks. They are not a function class specially designed to fit derivative estimates well. Similar to Neural ODEs, Liquid Neural Networks (liquidneuralnet) focus on the same class of problems and are similarly not applicable.
@@ B.3 Why didn’t you do experiments on FlyingThings3D / <simulated dataset>?
Most popular synthetic datasets do not contain long observation sequences (flyingthings; Butler:ECCV:2012), but instead include standalone frame pairs. Our method leverages the long sequence of observations to refine our neural estimate of the true ODE. Indeed, on two frames, EulerFlow collapses to NSFP.
More importantly, these datasets are also not representative of real world environments. To quote chodosh2023: “[FlyingThings3D has] unrealistic rates of dynamic motion, unrealistic correspondences, and unrealistic sampling patterns. As a result, progress on these benchmarks is misleading and may cause researchers to focus on the wrong problems.” khatri2024trackflow also make this point by highlighting the importance of meaningfully breaking down the object distribution during evaluation identify performance on rare safety-critical categories. FlyingThings3D does not have meaningful semantics; it’s not obvious what things even matter or how to appropriately break down the scene.
@@ D.1 Formulating the ODE
Given a (possibly moving) particle in some canonical frame (i.e. time 0), we define a function $L(x_{0},y_{0},z_{0},t)$ that can describe its location at an arbitrary future time $t$, i.e. a Lagrangian description of motion (Figure 4).
@@ D.2 Arbitrary start and end times from the Eulerian formulation
In the above derivation, $L$ requires that a moving point be defined in terms of a canonical frame defined at time 0, as is common in the deformation in reconstruction literature. However, the Eulerian formulation has no such requirement, allowing us to select arbitrary start and end times across different point queries. To showcase this, we can query $F$ to extract the trajectory of a particle at $t$ across the range $t,t^{\prime}]$ starting at $x_{t},y_{t},z_{t}$ simply by changing the range of the integral in Equation 10, i.e.
$E(x_{t},y_{t},z_{t},t,t^{\prime})=\left({x_{t},y_{t},z_{t}}\right)+\int_{t}^{t^{\prime}}F(x_{\tau},y_{\tau},z_{\tau},\tau)d\tau$ (11)
@@ D.3 Euler Integration to approximately solve the ODE
If $F$ is of arbitrary form and we want to compute the concrete values of $L$, we cannot exactly compute the continuous integral from 0 to $t$; we must approximate this with finite differences. Thus, we split the time range 0 to $t$ into $k$ steps, where each step is of size $\frac{t}{k}$. Thus, we can again define $L$ via recursion, but this time explicitly.